OpenStudy (anonymous):

which of the following is equal to the fraction below ? (4/5) ^6 a. 4^6 /5^6 b.6*(4/5) c.4^6 /5 d.24/30 ?? help

jimthompson5910 (jim_thompson5910):

Use the identity (x/y)^z = (x^z)/(y^z) to get (4/5)^6 = (4^6)/(5^6)

OpenStudy (anonymous):

a ?

jimthompson5910 (jim_thompson5910):

yes the answer is choice A